# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v2

@@ III-B3 Generator Enhancement
Technologies in prompt engineering [204] that focus on improving the quality of LLMs’ output, such as Prompt compression, Stepback Prompt [205], Active Prompt [206], Chain of Thought Prompt [174], etc., are all applicable to LLM generators in RAG systems. LLMLingua [207] applies a small model to compresses the overall length of the query to accelerate model inference, relieving the negative impact of irrelevant information on the model and alleviating the phenomenon of ”Lost in the Middle”[179]. ReMoDiffuse [50] decomposes complex descriptions into anatomical text scripts by using ChatGPT. ASAP [208] add exemplar tuples to the prompt for better results. An exemplar tuple is composed of the input code, a function definition, the results of analyzing that definition and its associated comment. CEDAR [131] uses a designed prompt template to organize code demonstration, query, and natural language instructions into a prompt. XRICL [194] utilizes COT technology to add translation pairs as an intermediate step in cross linguistic semantic parsing and inference. MakeAnAudio [43] is able to use other modalities as input which can provide much richer information for the following process.
@@
Decoding tuning refers to adding additional controls during the generator processing, which can be achieved by adjusting hyperparameters to achieve greater diversity, limiting the output vocabulary in some form, and so on. InferFix [132] balances the diversity and quality of results by adjusting the temperature in decoder. SYNCHROMESH [121] limits the output vocabulary of the decoder by implementing a completion engine to eliminate implementation errors.
Finetune Generator
The finetuning of the generator can enhance the model’s ability to have more precise domain knowledge or better fit with the retriever. RETRO [35] fixes the parameters of the retriever and uses the chunked cross attention mechanism in the generator to combine the content of the query and retriever. APICoder [130] finetunes the generator CODEGEN-MONO 350M [209] with a shuffled new file combined with API information and code blocks. CARE [159] first uses image data, audio data and vedio-text pairs to train encoders and then finetune the decoder (generator) with the target of decreasing caption loss and concept detection loss together, during which the encoders and the retriever are frozen. Animate-A-Story [210] optimizes the video generator with image data, and then finetunes a LoRA [211] adapter to capture the appearance details of the given character. RetDream [49] finetunes a LoRA adapter [211] with the rendered images.
@@ III-B4 Result Enhancement
In many scenarios, the final result of RAG may not achieve the expected effect, and some techniques of Result Enhancement can help alleviate this problem.
Rewrite Output
SARGAM [212] modifies the generated results in code-related task by classifying Deletion Classifier, Placeholder Classifier and Insertion Classifier with Levenshtein Transformer to better adapt to the actual code context. Ring [213] obtains diversity results by reranking candidates based on the average of per token log probabilities produced by the generator. CBR-KBQA [53] revises the result by aligning generated relations with those presented in the local neighborhood of the query entity in knowledge graph.
@@ III-B5 RAG Pipeline Enhancement
Rule-based: FLARE [214] actively decides whether and when to search through the probability in the generation process. Efficient-KNNLM [37] combines the generation probability of KNN-LM [36] and NPM [161] with a hyperparameter $\lambda$ to determine the proportion of generation and retrieval. Mallen et al. [215] conducts statistical analysis on questions before generating answers, allowing the model to directly answer high-frequent questions, and introducing RAG for low-frequent questions. Jiang et al. [216] studies Model Uncertainty, Input Uncertainty, and Input Statistics to comprehensively assess the confidence level of the model. Ultimately, based on the confidence level of the model, a decision is made whether to retrieve. Kandpal et al. [217] assists in determining whether a retrieval is needed by studying the relationship between the number of relevant documents in the training dataset and the degree to which the model masters relevant knowledge.
Model-based: Self-RAG [127] uses a trained generator to determine whether to perform a retrieval based on the retrieve token under different instructions, and evaluates the relevance and level of support of the retrieved text through the Self-Reflection token. Finally, the quality of the final output result is evaluated based on the Critique token. Ren et al. [218] uses ”Judgment Prompting” to determine whether LLMs can answer relevant questions and whether their answers are correct or not, thereby assisting in determining the necessity of a retrieval. SKR [219] uses the ability of LLMs themselves to judge in advance whether they can answer the question, and if they can answer, no retrieval is performed. Rowen [220] employs a model as a sophisticated multilingual detection system to evaluate the semantic coherence of answers to identical questions posed across various languages. In the event of detected inconsistencies, it decides to retrieve external information, thereby enhancing the reasoning process and rectifying inaccuracies. Conversely, when responses exhibit consistency, the system upholds the initially generated answer, which is derived from internal reasoning. AdaptiveRAG [221] dynamically decides whether to retrieve based on the query complexity by a classifier, which is a smaller LM.
@@
RepoCoder [222] applies an iterative retrieval-generation pipeline to better utilize the useful information scattered in different files in the code completion tasks. It augments the retrieval query during the i-th iteration with previously generated code and obtain greater results. ITER-RETGEN [223] synergizes retrieval and generation in an iterative manner. The current output of the generator can to some extent reflect the knowledge it still lacks, and the retrieve can retrieve the missing information as contextual information for the next round, which helps to improve the quality of the generated content in the next round. SelfMemory [224] employs a retrieval-augmented generator in an iterative manner to create an unlimited memory pool. Following this, a memory selector is used to choose one output, which then serves as the memory for the subsequent generation round.
@@ IV Applications
Figure 8: Taxonomy of multiple applications in multimodality.
In this chapter, we will introduce the popular RAG applications in different fields.
@@ IV-A RAG for Text
In this section, we will introduce some popular work for text generation using RAG methods.
@@ IV-A1 Question Answering
FID [34] and REALM [32] identify the top k most pertinent article snippets based on the query and forward each snippet along with the question to LLMs to generate k responses. These responses are then synthesized into a final answer. Toutanova et al. [225] substituted the text corpus in REALM with subgraphs from a knowledge graph, yielding impressive results. Atlas [29] demonstrated that leveraging RAG to assist LLMs in open-domain question answering tasks significantly enhances few-shot learning capabilities. As shown in Fig. 9, RETRO [35] employs attention mechanisms to integrate the question with relevant retrieved documents within the model to produce the final answer. SKR [219] observed that using RAG does not invariably benefit Question Answering and thus explored guiding the model to evaluate its grasp of pertinent knowledge, subsequently adapting its use of external resources for retrieval enhancement. TOG [226] introduced an innovative knowledge graph-augmented LLM framework, which excels by fostering interactions between LLMs and the Knowledge Graph and by expanding the inference path space with beam search. NPM [161] pioneered the use of nonparametric data distributions in lieu of the softmax layer, enabling models with fewer parameters to perform effectively. Self-RAG [127] improves answer quality by learning to discern when to retrieve, assess the retrieved content’s relevance, and evaluate the final generated results using four types of reflective tokens. CL-ReLKT [227] employs a language-generalized encoder to bridge the gap between question-document pairs across languages, thus better leveraging multilingual data. CORE [228] mitigates language resource disparities by introducing a novel dense passage retrieval algorithm and a multilingual autoregressive generation model. Lastly, EAE [154] enhances answer quality by retrieving entity embeddings for query entities and integrating these with hidden states for further processing.
@@
UR-QA [229] found that when encountering unseen problems, retrieving QA pairs has a better final effect; When encountering problems that have not been seen before, the retrieve text chunk performs better. Therefore, it is proposed to simultaneously retrieve QA pairs and text chunks, and select the final answer by comparing the calibrated confidences. ARM-RAG [230] demonstrates exceptional performance in solving elementary mathematics problems by retrieving past successful reasoning steps. DISC-LawLLM [231] constructs a supervised fine-tuning dataset through a legal syllogism prompting strategy, enabling the model to receive support from the latest legal information. RAG-end2end [232] conducts simultaneous training of the retriever(DPR) and the generator(BART) to optimize performance for the end-to-end question-answering task and to facilitate domain adaptation. MultiHop-RAG [233] is designed to extract pertinent information from a variety of distinct documents, aggregating this knowledge to equip the generator with the necessary context for producing the definitive answer to the query.
@@ IV-A2 Fact Verification
Fact Verification involves assessing the veracity of information, a critical function in disciplines such as Natural Language Processing (NLP), Information Retrieval, and Data Mining. In today’s digital era, characterized by an exponential increase in data, particularly across social media and online news platforms, there is a rapid proliferation of unchecked information. Fact verification plays an essential role in countering the spread of fake news, deceptive content, and rumors, thereby preserving the integrity of the information landscape and ensuring the public’s access to accurate knowledge. Consequently, automated fact verification systems are of immense importance, with broad applications and significant practical value.
CONCRETE [234] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Hagström et al. [235] proves on LLaMA [236] and Atlas [29] that search augmentation is more beneficial for solving inconsistency problems than increasing model size.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. However, the vast scope of common sense knowledge and the intricacies of reasoning processes make Commonsense Reasoning a perennially challenging and prominent area of research within the field of NLP.
KG-BART [237] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. This approach not only improves the models’ generalization capabilities but also significantly bolsters their Commonsense Reasoning proficiency. Wan et al. [238] constructs the CONFLICTINGQA dataset, comprising contentious questions and conflicting answer documents, to examine which textual features significantly influence LMs’ ability to independently navigate controversial issues. The findings reveal that LMs often neglect the stylistic aspects of text that are typically valued by people.
@@ IV-A4 Human-Machine Conversation
Human-Machine Conversation encompasses the ability of machines to comprehend natural language and adeptly employ this skill to engage with humans seamlessly. This capability represents a significant challenge within the realms of Artificial Intelligence and Natural Language Processing and offers a broad spectrum of practical applications. As such, Human-Machine Conversation continues to be a focal point of research for many scholars.
ConceptFlow [239] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. This method achieves commendable results even with a substantial reduction in model parameters. Cai et al. [240] reimagines the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [241] augments dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [242] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [243], PARC [244], and CREA-ICL [245] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [246] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval. If the answer is correct, the retrieved document is added to the original answer as a reference; if the answer lacks reliable references, it guides the LLM to respond to the question anew.

### v3

@@ III-B3 Generator Enhancement
Technologies in prompt engineering [199] that focus on improving the quality of LLMs’ output, such as Prompt compression, Stepback Prompt [200], Active Prompt [201], Chain of Thought Prompt [173], etc., are all applicable to LLM generators in RAG systems. LLMLingua [202] applies a small model to compresses the overall length of the query to accelerate model inference, relieving the negative impact of irrelevant information on the model and alleviating the phenomenon of “Lost in the Middle”[155]. ReMoDiffuse [51] decomposes complex descriptions into anatomical text scripts by using ChatGPT. ASAP [203] add exemplar tuples to the prompt for better results. An exemplar tuple is composed of the input code, a function definition, the results of analyzing that definition and its associated comment. CEDAR [130] uses a designed prompt template to organize code demonstration, query, and natural language instructions into a prompt. XRICL [189] utilizes COT technology to add translation pairs as an intermediate step in cross linguistic semantic parsing and inference. ACTIVERAG [204] employs the Cognition Nexus mechanism to calibrate the intrinsic cognition of LLMs and applies COT prompt in answer generation. Make-An-Audio [44] is able to use other modalities as input which can provide much richer information for the following process.
@@
Decoding tuning refers to adding additional controls during the generator processing, which can be achieved by adjusting hyperparameters to achieve greater diversity, limiting the output vocabulary in some form, and so on.
InferFix [131] balances the diversity and quality of results by adjusting the temperature in decoder. SYNCHROMESH [122] limits the output vocabulary of the decoder by implementing a completion engine to eliminate implementation errors.
Generator Finetuning
The finetuning of the generator can enhance the model’s ability to have more precise domain knowledge or better fit with the retriever.
RETRO [36] fixes the parameters of the retriever and uses the chunked cross attention mechanism in the generator to combine the content of the query and retriever. APICoder [129] finetunes the generator CODEGEN-MONO 350M [205] with a shuffled new file combined with API information and code blocks. CARE [160] first uses image data, audio data and vedio-text pairs to train encoders and then finetune the decoder (generator) with the target of decreasing caption loss and concept detection loss together, during which the encoders and the retriever are frozen. Animate-A-Story [206] optimizes the video generator with image data, and then finetunes a LoRA [207] adapter to capture the appearance details of the given character. RetDream [50] finetunes a LoRA adapter [207] with the rendered images.
@@ III-B4 Result Enhancement
In many scenarios, the result of RAG may not achieve the expected effect, and some techniques of Result Enhancement can help alleviate this problem.
Output Rewrite
Output Rewrite refers to rewriting the content generated by the generator in certain scenarios to meet the needs of downstream tasks.
SARGAM [208] refines outputs in code-related tasks by employing a special Transformer alongside Deletion, Placeholder, and Insertion Classifiers to better align with the real-world code context. Ring [209] obtains diversity results by reranking candidates based on the average of per token log probabilities produced by the generator. CBR-KBQA [54] revises the result by aligning generated relations with those presented in the local neighborhood of the query entity in knowledge graph.
@@ III-B5 RAG Pipeline Enhancement
RAG Pipeline Enhancement refers to optimizing the processes of RAG at the system level in order to achieve better performance results.
@@
Rule-based: FLARE [210] actively decides whether and when to search through the probability in the generation process. Efficient-KNNLM [38] combines the generation probability of KNN-LM [37] and NPM [162] with a hyperparameter $\lambda$ to determine the proportion of generation and retrieval. Mallen et al. [211] used statistical analysis on questions to enable direct answers for high-frequency ones and applied RAG for low-frequency ones. Jiang et al. [212] studied Model Uncertainty, Input Uncertainty, and Input Statistics to comprehensively assess the confidence level of the model. Ultimately, based on the confidence level of the model, a decision is made whether to retrieve. Kandpal et al. [213] studied the correlation between the number of relevant documents and the model’s knowledge mastery to assess the need for retrieval.
Model-based: Self-RAG [126] uses a trained generator to determine whether to perform a retrieval based on the retrieve token under different instructions, and evaluates the relevance and level of support of the retrieved text through the Self-Reflection token. Finally, the quality of the final output result is evaluated based on the Critique token. Ren et al. [214] used ”Judgment Prompting” to determine whether LLMs can answer relevant questions and whether their answers are correct or not, thereby assisting in determining the necessity of a retrieval. SKR [215] uses the ability of LLMs themselves to judge in advance whether they can answer the question, and if they can answer, no retrieval is performed. Rowen [216] employs a model as a sophisticated multilingual detection system to evaluate the semantic coherence of answers to identical questions posed across various languages. In the event of detected inconsistencies, it decides to retrieve external information, thereby enhancing the reasoning process and rectifying inaccuracies. Conversely, when responses exhibit consistency, the system upholds the initially generated answer, which is derived from internal reasoning. AdaptiveRAG [217] dynamically decides whether to retrieve based on the query complexity by a classifier, which is a smaller LM.
@@
Iterative RAG progressively refines results by repeatedly cycling through retrieval and generation phases, rather than a single round.
RepoCoder [218] employs an iterative retrieval-generation pipeline for code completion tasks, enhancing each retrieval query with code generated in prior iterations to more effectively leverage information dispersed across various files, thereby achieving superior results. ITER-RETGEN [219] synergizes retrieval and generation in an iterative manner. The current output of the generator can to some extent reflect the knowledge it still lacks, and the retrieve can retrieve the missing information as contextual information for the next round, which helps to improve the quality of the generated content in the next round. SelfMemory [220] employs a retrieval-augmented generator in an iterative manner to create an unlimited memory pool. Following this, a memory selector is used to choose one output, which then serves as the memory for the subsequent generation round. RAT [221] initial generates content by an LLM with a zero-shot CoT prompt, then revises each thought step by retrieving knowledge from external knowledge base.
@@ IV Applications
TABLE I: Taxonomy of RAG applications across various modalities.
RAG for Text
Question Answering Human-Machine Conversation Neural Machine Translation Summarization Others
REALM§ TKEGEN§ TOG SKR§ Self-RAG§ RIAG FiD§ RETRO§ NPM§ CREA-ICL BlenderBot3§ CEG Internet-Augmented-DG§ ConceptFlow§ Skeleton-to-Response§ NMT-with-Monolingual-TM§ TRIME§ KNN-MT§ COG RAMKG§ RPRR RIGHT§ Unlimiformer§ CONCRETE§ Atlas§ KG-BART§ R-GQA§
RAG for Code
Code Generation Code Summarization Code Completion Automatic Program Repair Text-to-SQL and Code -based Semantic Parsing Others
SKCODER§ RRGCode ARKS RECODE KNN-TRANX Toolcoder§ RACE BASHEXPLAINER READSUM Rencos CoRec Tram§ EDITSUM ReACC RepoCoder§ De-Hallucinator REPOFUSE§ RepoFusion§ EDITAS§ RING CEDAR§ RAP-Gen§ InferFix§ SARGAM§ RTLFixer§ XRICL§ SYNCHROMESH§ RESDSQL§ REFSQL§ CodeICL§ MURRE De-fine Code4UIE§ E&V StackSpotAI§ ImputBlaster
RAG for Knowledge RAG for 3D
Knowledge Base QA Knowledge-augmented Open-domain QA Table for QA Others Text-to-3D
CBR-KBQA§ TIARA§ Keqing§ RNG-KBQA ReTraCk§ SKP§ UniK-QA KG-FiD GRAPE SKURG KnowledGPT EFSUM§ EfficientQA CORE§ Convinse RINK§ T-RAG§ StructGPT GRetriever§ SURGE§ K-LaMP RHO ReMoDiffuse AMD
RAG for Image RAG for Video
Image Generation Image Captioning Others Video Captioning Video QA & Dialogue Others
RetrieveGAN IC-GAN§ Re-imagen§ RDM Retrieve&Fuse§ KNN-Diffusion MA REVEAL SMALLCAP CRSR RA-Transformer PICa Maira KIF RA-VQA KaVD§ R-ConvED§ CARE§ EgoInstructor§ MA-DRNN R2A Tvqa+§ VGNMN VidIL RAG-Driver Animate-A-Story§
RAG for Science RAG for Audio
Drug Discovery Biomedical Informatics Enhancement Math Applications Audio Generation Audio Captioning
RetMol§ PromptDiff PoET Chat-Orthopedist§ BIOREADER MedWriter QARAG LeanDojo RAG-for-math-QA Re-AudioLDM§ Make-An-Audio§ RECAP§
Query-based
Latent-based
Logit-based
Speculative
Query+LatentLatent+Logit
[…]
In this section, we focus on RAG applications spanning various modalities. To echo with the taxonomy of RAG foundations and enhancements, we also demonstrate their utilization across different tasks in Table I.
@@ IV-A RAG for Text
To begin with, text generation is among the most important and widely deployed applications for RAG. Here we introduce popular works for seven tasks, respectively.
@@ IV-A1 Question Answering
FiD [35] and REALM [33] identify the top-k most pertinent article snippets based on the query and forward each snippet along with the question to LLMs to generate k responses. These responses are then synthesized into a final answer. Toutanova et al. [222] substituted the text corpus in REALM with subgraphs from a knowledge graph, yielding impressive results. As shown in Fig. 7, RETRO [36] employs attention mechanisms to integrate the question with relevant retrieved documents within the model to produce the final answer. SKR [215] observes that using RAG does not invariably benefit Question Answering and thus explored guiding the model to evaluate its grasp of pertinent knowledge, subsequently adapting its use of external resources for retrieval enhancement. TOG [223] introduces an innovative knowledge graph-augmented LLM framework, which excels by fostering interactions between LLMs and the Knowledge Graph and by expanding the inference path space with beam search. NPM [162] pioneers the use of nonparametric data distributions in lieu of the softmax layer, enabling models with fewer parameters to perform effectively. Self-RAG [126] improves answer quality by learning to discern when to retrieve, assess the retrieved content’s relevance, and evaluate the final generated results using four types of reflective tokens. CL-ReLKT [224] employs a language-generalized encoder to bridge the gap between question-document pairs across languages, thus better leveraging multilingual data. CORE [225] mitigates language resource disparities by introducing a novel dense passage retrieval algorithm and a multilingual autoregressive generation model. Lastly, EAE [156] enhances answer quality by retrieving entity embeddings for query entities and integrating these with hidden states for further processing.
@@
UR-QA [226] found that when encountering unseen problems, retrieving QA pairs has a better final effect; When encountering problems that have not been seen before, the retrieve text chunk performs better. Therefore, it is proposed to simultaneously retrieve QA pairs and text chunks, and select the final answer by comparing the calibrated confidences. DISC-LawLLM [227] constructs a supervised fine-tuning dataset through a legal syllogism prompting strategy, enabling the model to receive support from the latest legal information. RAG-end2end [228] conducts simultaneous training of the retriever (DPR) and the generator (BART) to optimize performance for the end-to-end question-answering task and to facilitate domain adaptation. MultiHop-RAG [229] is designed to extract pertinent information from a variety of distinct documents, aggregating this knowledge to equip the generator with the necessary context for producing the definitive answer to the query.
@@ IV-A2 Fact Verification
Fact Verification involves assessing the veracity of information, a critical function in disciplines such as Natural Language Processing (NLP), Information Retrieval, and Data Mining. In today’s digital era, characterized by an exponential increase in data, particularly across social media and online news platforms, there is a rapid proliferation of unchecked information. Fact verification plays an essential role in countering the spread of fake news, deceptive content, and rumors, thereby preserving the integrity of the information landscape and ensuring the public’s access to accurate knowledge. Consequently, automated fact verification systems are of immense importance, with broad applications and significant practical value. CONCRETE [230] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Hagström et al. [231] proved on LLaMA [4] and Atlas [30] that search augmentation is more beneficial for solving inconsistency problems than increasing model size. Atlas [30] shows that using RAG to support LLMs in knowledge-intensive tasks markedly improves their few-shot learning performance.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. However, the vast scope of common sense knowledge and the intricacies of reasoning processes make Commonsense Reasoning a perennially challenging and prominent area of research within the field of NLP. KG-BART [232] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. This approach not only improves the models’ generalization capabilities but also significantly bolsters their Commonsense Reasoning proficiency. Wan et al. [233] constructed the CONFLICTINGQA dataset, comprising contentious questions and conflicting answer documents, to examine which textual features significantly influence LMs’ ability to independently navigate controversial issues. The findings reveal that LMs often neglect the stylistic aspects of text that are typically valued by people.
@@ IV-A4 Human-Machine Conversation
Human-Machine Conversation encompasses the ability of machines to comprehend natural language and adeptly employ this skill to engage with humans seamlessly. This capability represents a significant challenge within the realms of Artificial Intelligence and Natural Language Processing and offers a broad spectrum of practical applications. As such, Human-Machine Conversation continues to be a focal point of research for many scholars. ConceptFlow [234] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. This method achieves commendable results even with a substantial reduction in model parameters. Cai et al. [235] reimagined the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [236] augmented dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [237] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [238], PARC [239], and CREA-ICL [240] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [241] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval. If the answer is correct, the retrieved document is added to the original answer as a reference; if the answer lacks reliable references, it guides the LLM to respond to the question anew.